OpenStudy (anonymous):

2 cards are drawn without replacement from a standard deck of 52 cards. What is the probability that both are spades given that one is a spade.

OpenStudy (anonymous):

So how many spades are in a deck of cards?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

What is the probability of pulling a spade for the first card?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

okay... lets keep the fractions in their original numbers... 13/52. this will make more sense. Now, the second card, you have pulled out the one spade already... now how many spades do you have left?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

good. If we're not going to replace that card back in to the deck... how many total cards do we have left in the deck?

OpenStudy (anonymous):

51

OpenStudy (anonymous):

good. so what is the probability of pulling out a spade the second time?

OpenStudy (anonymous):

12/51

OpenStudy (anonymous):

now.. with two card pulls, stats calls it an event... based on the word problem.. you need to pull one card and then another card... when you see the word... and it means you need to multiple the two events or ratios together.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

the question asks "What is the probability that both are spades given that one is a spade."

OpenStudy (anonymous):

the "given" part means you KNOW the first card was a spade so there are 12 spades, 51 in the deck total, and the probability that you draw a spade is now \(\frac{12}{51}\) and that is the answer to the question

OpenStudy (anonymous):

Um you don't know that the first one it the spade. The second card could also be the spade draw.

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

right.. the probability of pulling the spades... the problem isnt asking whats the probability of not pulling a spade.. that's a different problem.

OpenStudy (anonymous):

yes, you do know it that is what "given" means in "2 cards are drawn without replacement from a standard deck of 52 cards. What is the probability that both are spades GIVEN that one is a spade"

OpenStudy (anonymous):

ooh i see, it is "given one is a spade" not "given the first one is a spade"

OpenStudy (anonymous):

Yes and also the question just before this one asked the probability given that the first draw was a spade so I don't think I would be asked the same question twice.

OpenStudy (anonymous):

The answer is 13.3% btw. I just don't how to get that answer.

OpenStudy (anonymous):

ok we can do this the probability that both are spades is \[\frac{13}{52}\times \frac{12}{51}\] the probability that at least one of the two cards is a space is \[1-\frac{49}{52}\times \frac{48}{51}\]

OpenStudy (anonymous):

and so since \[P(A|B)=\frac{P(A\cap B)}{P(B)}\] you get \[\frac{\frac{13}{52}\times \frac{12}{51}}{1-\frac{49}{52}\times \frac{48}{51}}\]

OpenStudy (anonymous):

there you go... after you multiple the two.. the GIVEN part of the question... is the next step

OpenStudy (anonymous):

13% exactly which makes me wonder if there is an easier way to do it http://www.wolframalpha.com/input/?i=%2813%2F52-12%2F51%29%2F%281-49%2F52*48%2F51%29

OpenStudy (anonymous):

@kealiiballao yeah, i should learn to read is there an easier way?

OpenStudy (anonymous):

I really have no other way doing this types of problems...

OpenStudy (anonymous):

@ArkGoLucky does that last step make sense to you?

OpenStudy (anonymous):

1 = the probably of pulling any card 49/52 = on the first event, the probability of not pulling an ace. 48/51 = on the second event, the probably not pulling an ace the second time.

OpenStudy (anonymous):

I understand the formula but would the probability of getting at least one spade be 1-(39/52)*(38/51)

OpenStudy (anonymous):

spades not aces

OpenStudy (anonymous):

oh yeah.. spades

OpenStudy (anonymous):

Nope... because the bottom part.. is based on the opposite situation.

OpenStudy (anonymous):

we could try this \[\frac{\frac{13}{52}\times \frac{12}{51}}{2\times \frac{13}{52}\times \frac{49}{51}+\frac{13}{52}\times \frac{12}{52}}\]

OpenStudy (anonymous):

numerator is both are spades denominator is 1st is, second isn't or 1st isn't second is or both are

OpenStudy (anonymous):

then this gives \[\frac{13\times 12}{2\times 13\times 49+12\times 13}\]

OpenStudy (anonymous):

hmm this isn't 13% though...

OpenStudy (anonymous):

oh yes it is! i made a typo!!

OpenStudy (anonymous):

i put 49 instead of 39 doh now it is right http://www.wolframalpha.com/input/?i=13*12%2F%282*13*39%2B13*12%29

OpenStudy (anonymous):

did it twice!!\[\frac{\frac{13}{52}\times \frac{12}{51}}{1-\frac{39}{52}\times \frac{38}{51}}\]

OpenStudy (anonymous):

also here \[\frac{\frac{13}{52}\times \frac{12}{51}}{2\times \frac{13}{52}\times \frac{39}{51}+\frac{13}{52}\times \frac{12}{52}}\]

OpenStudy (anonymous):

now it is correct, you get \(13.333\%\)

OpenStudy (anonymous):

How did you know that P(B) was first card, second card or both?

OpenStudy (anonymous):

B is "one card is a spade" so it means one or the other or both

OpenStudy (anonymous):

i.e. first or second or both first and second

OpenStudy (anonymous):

Okay I think I understand.

OpenStudy (anonymous):

let me check that both solutions give the same answer

OpenStudy (anonymous):

yeah they both give the same http://www.wolframalpha.com/input/?i=%2813%2F52*12%2F51%29%2F%281-39%2F52*38%2F51%29